OpenStudy (anonymous):

PLEASE SIMPLIFY AND SHOW STEPS: ((2/X)-3)/(1-(1/(X-1))) THANK YOU!!!!!!!!!!!!!

OpenStudy (anonymous):

first you need to make common denominators so (2/x)-3=2/x - 3x/x=-1/x 1-1/(x-1)=(x-1)/(x-1) -1/(x-1)=(x-2)/(x-1) Now dividing is the same as multiplying by the reciprocal (so the 1/x) So (-1/x) / (x-2)/(x-1)=(-1/x) *(x-1)/(x-2)=(1-x)/(x^2-2x)

OpenStudy (anonymous):

Might have done a minor mistake because the result is not the simple. But I guess you have the idea.

OpenStudy (anonymous):

I see (2/x)-3=2/x - 3x/x=(2-3x)/x So it will be ((2-3x)/x)/((x-2)/(x-1))=((2-3x)/x)*((x-1)/(x-2)) =((2-3x)(x-1))/(x*(x-2)=(2x-2-3x^2+3x)/((x^2-2x) =(-3x^2+5x-2)/(x^2-2x)